# Thermal Hall transport in Kitaev spin liquids

### v1

@@ IV.2.2 Symmetric off-diagonal interaction $\Gamma$
Next, we investigate the effects of the $\Gamma$ interaction. Here, we consider two representative magnetic fields, $h=0.04$ (Fig. 10) and $h=0.08$ (Fig. 11). Figures 10(a)–10(c) and 11(a)–11(c) show the temperature dependence of the specific heat, the magnetization, and the flux density for $-0.02\leq\Gamma\leq 0.02$. Although the magnitudes of $\Gamma$ are relatively small, we find that the low-temperature peak of the specific heat undergoes significant changes: it shifts to higher (lower) temperatures and its height increases (decreases) for negative (positive) $\Gamma$. At $h=0.08$, the peak nearby disappears, and the specific heat exhibits less temperature dependence at low temperatures at $\Gamma=0.02$. Similarly, the magnetization increases (decreases) for negative (positive) $\Gamma$. We also observe $\Gamma$-sign-dependent changes in the flux density around the temperature region $T\sim 0.03$, where the specific heat exhibits the low-temperature peak, although the effect is less pronounced.
@@ IV.3 Summary of $h$-$T$ dependence
Summarizing results in the previous subsections, we plot the temperature and field dependence of the specific heat, magnetization, flux density, and $\kappa_{xy}/T$ in Figs. 12, 13, 14, and 15, respectively, for various values of $\Gamma$ and $\Gamma^{\prime}$. In each figure, the central panel corresponds to the result for the pure Kitaev model with $K=-1$ and $\Gamma=\Gamma^{\prime}=0$. The upper panels display the results for varying $\Gamma^{\prime}$ with $\Gamma=0$, and the lower panels show the results for varying $\Gamma$ with $\Gamma^{\prime}=0$.
@@ IV.4 Classical limit
Figure 17: Corresponding plots to Fig. 16 under the magnetic field parallel to $[111]$, $[11\bar{2}]$, and $[\bar{1}10]$ with $h=0.04$.
Next, we discuss the field-direction dependence of physical quantities in the pure Kitaev model. Following the quantum case in Sec. IV.1.2, we consider three mutually-perpendicular field directions: $[11\bar{2}]$, $[\bar{1}10]$, and $[111]$. Figures 17(a) and 17(b) show the temperature dependence of the specific heat and magnetization, respectively. These results indicate that the field direction has little effect on either quantity, similar to the quantum system (see Fig. 8). In contrast to these bulk physical quantities, the thermal Hall conductivity strongly depends on the direction of the applied magnetic field. As shown in Fig. 17(c), $\kappa_{xy}$ is positive and increases with decreasing temperature for $\bm{h}\parallel[111]$, whereas it is negative and decreases for $\bm{h}\parallel[11\bar{2}]$; it vanishes for the $\bm{h}\parallel[\bar{1}10]$, consistent with the symmetry argument in the end of Sec. IV.1.2. Moreover, we observe that the absolute value for $\bm{h}\parallel[111]$ is larger than that for $\bm{h}\parallel[11\bar{2}]$. While these trends are consistent with those observed in the quantum system, the overshooting behavior with a peak structure in Fig. 8(d) is not reproduced in the classical case. Furthermore, although the three-spin contribution $\kappa_{xy}^{(L)}$, shown in Fig. 17(d), follows these trends, the two-spin contribution $\kappa_{xy}^{(M)}$, shown in Fig. 17(e), remains small and fails to reproduce the quantum behavior presented in Fig. 8(f). This result supports the significant role of quantum fluctuations in the two-spin contribution to the thermal Hall effect.
@@ V Summary
The enhanced $\kappa_{xy}/T$ with overshooting behavior is observed across a broad range of magnetic fields and interaction parameters. Moreover, the peak temperature of $\kappa_{xy}/T$ consistently aligns with the low-temperature peak of the specific heat, underscoring their common origin in fractional excitations. Our comprehensive analyses consistently support the dominant contributions of topological Majorana fermions to this overshooting behavior, rather than topological magnons. We note that previous studies based on the topological magnon picture predicted an opposite sign of $\kappa_{xy}/T$ for the pure Kitaev model [42, 94], although the inclusion of additional $\Gamma$ and $\Gamma^{\prime}$ interactions may alter the sign [40, 41]. One possible origin of this discrepancy lies in the limited applicability of the magnon picture in this regime, where the magnetization remains small. Indeed, it was demonstrated that magnon-magnon interactions beyond the linear spin-wave theory can alter not only the amplitude but also the sign of $\kappa_{xy}/T$ in this field and temperature regime [94]. Another plausible explanation involves how the edges were treated in previous calculations on topological magnons, which relied on the bulk-edge correspondence without explicitly accounting for edge-specific effects, such as magnetization reconstructions near the edges [97, 98].
@@ Appendix B Benchmark on thermal pure quantum state
In this Appendix, we examine the accuracy of the cTPQ state method by comparing its results with those obtained from the ED for a small $(L,L^{\prime})=(2,4)$ cluster (the total system size is $N_{\rm s}=16$). This serves as a preliminary step for the comparison between the cTPQ state and XTRG methods presented in Appendix C. In the ED approach, we diagonalize the Hamiltonian, whose dimension is $655362^{16}=$ 65536, using ScaLAPACK [100], and compute the temperature dependence of physical quantities based on the obtained eigenvalues and eigenvectors. In the cTPQ state method, we employ the calculation conditions mentioned in Sec. III.3.

### v2

@@ IV.2.2 Symmetric off-diagonal interaction $\Gamma$
Next, we investigate the effects of the $\Gamma$ interaction. Here, we consider two representative magnetic fields, $h=0.04$ (Fig. 10) and $h=0.08$ (Fig. 11). Figures 10(a)–10(c) and 11(a)–11(c) show the temperature dependence of the specific heat, the magnetization, and the flux density for $0.02\leq\Gamma\leq 0.02$. Although the magnitudes of $\Gamma$ are relatively small, we find that the low-temperature peak of the specific heat undergoes significant changes: it shifts to higher (lower) temperatures and its height increases (decreases) for negative (positive) $\Gamma$. At $h=0.08$, the peak nearby disappears, and the specific heat exhibits less temperature dependence at low temperatures at $\Gamma=0.02$. Similarly, the magnetization increases (decreases) for negative (positive) $\Gamma$. We also observe $\Gamma$-sign-dependent changes in the flux density around the temperature region $T\sim 0.03$, where the specific heat exhibits the low-temperature peak, although the effect is less pronounced.
@@ IV.3 Summary of $h$-$T$ dependence
Summarizing results in the previous subsections, we plot the temperature and field dependence of the specific heat, magnetization, flux density, and $\kappa_{xy}/T$ in Figs. 12, 13, 14, and 15, respectively, for various values of $\Gamma$ and $\Gamma^{\prime}$. In each figure, the central panel corresponds to the result for the pure Kitaev model with K=1K=-1 and $\Gamma=\Gamma^{\prime}=0$. The upper panels display the results for varying $\Gamma^{\prime}$ with $\Gamma=0$, and the lower panels show the results for varying $\Gamma$ with $\Gamma^{\prime}=0$.
@@ IV.4 Classical limit
Figure 17: Corresponding plots to Fig. 16 under the magnetic field parallel to $[111]$, $11\bar{2}]$, and $\bar{1}10]$ with $h=0.04$.
Next, we discuss the field-direction dependence of physical quantities in the pure Kitaev model. Following the quantum case in Sec. IV.1.2, we consider three mutually-perpendicular field directions: $1\bar{2}], [1¯1$0][$\bar${1}10], and $[111]$. Figures 17(a) and 17(b) show the temperature dependence of the specific heat and magnetization, respectively. These results indicate that the field direction has little effect on either quantity, similar to the quantum system (see Fig. 8). In contrast to these bulk physical quantities, the thermal Hall conductivity strongly depends on the direction of the applied magnetic field. As shown in Fig. 17(c), $\kappa_{xy}$ is positive and increases with decreasing temperature for $\bm{h}\parallel[111]$, whereas it is negative and decreases for $\bm{h}\parallel[11\bar{2}]$; it vanishes for the $\bm{h}\parallel[\bar{1}10]$, consistent with the symmetry argument in the end of Sec. IV.1.2. Moreover, we observe that the absolute value for $\bm{h}\parallel[111]$ is larger than that for $\bm{h}\parallel[11\bar{2}]$. While these trends are consistent with those observed in the quantum system, the overshooting behavior with a peak structure in Fig. 8(d) is not reproduced in the classical case. Furthermore, although the three-spin contribution $\kappa_{xy}^{(L)}$, shown in Fig. 17(d), follows these trends, the two-spin contribution $\kappa_{xy}^{(M)}$, shown in Fig. 17(e), remains small and fails to reproduce the quantum behavior presented in Fig. 8(f). This result supports the significant role of quantum fluctuations in the two-spin contribution to the thermal Hall effect.
@@ V Summary
The enhanced $\kappa_{xy}/T$ with overshooting behavior is observed across a broad range of magnetic fields and interaction parameters. Moreover, the peak temperature of $\kappa_{xy}/T$ consistently aligns with the low-temperature peak of the specific heat, underscoring their common origin in fractional excitations. Our comprehensive analyses consistently support the dominant contributions of topological Majorana fermions to this overshooting behavior, rather than topological magnons. We note that previous studies based on the topological magnon picture predicted an opposite sign of $\kappa_{xy}/T$ for the pure Kitaev model [42, 95], although the inclusion of additional $\Gamma$ and $\Gamma^{\prime}$ interactions may alter the sign [40, 41]. One possible origin of this discrepancy lies in the limited applicability of the magnon picture in this regime, where the magnetization remains small. Indeed, it was demonstrated that magnon-magnon interactions beyond the linear spin-wave theory can alter not only the amplitude but also the sign of $\kappa_{xy}/T$ in this field and temperature regime [95]. Another plausible explanation involves how the edges were treated in previous calculations on topological magnons, which relied on the bulk-edge correspondence without explicitly accounting for edge-specific effects, such as modulation of the magnetic states near the edges [98, 99]. Indeed, our calculations reveal that the edge magnetization differs significantly from the bulk magnetization as discussed in Appendix E.
@@ Appendix B Benchmark on thermal pure quantum state
In this Appendix, we examine the accuracy of the cTPQ state method by comparing its results with those obtained from the ED for a small $(L,L^{\prime})=(2,4)$ cluster (the total system size is $N_{\rm s}=16$). This serves as a preliminary step for the comparison between the cTPQ state and XTRG methods presented in Appendix C. In the ED approach, we diagonalize the Hamiltonian, whose dimension is $2^{16}=65536$, using ScaLAPACK [101], and compute the temperature dependence of physical quantities based on the obtained eigenvalues and eigenvectors. In the cTPQ state method, we employ the calculation conditions mentioned in Sec. III.3.
@@ Appendix C Comparison between thermal pure quantum state method and tensor network method
Appendix D Spatial extension of the energy current
In this Appendix, we present the supplemental data for $l$ dependence of the energy current amplitude for the pure Kitaev model. As discussed in Sec. III.1, our definition of the thermal Hall conductivity $\kappa_{xy}$ relies on the assumption that the energy current is well localized at the edge. In Fig. 7, we showed that the energy current amplitude is actually well localized at the edge for $T\simeq 0.0186$. An almost exponential decay of the energy current amplitude is also observed at higher temperatures, as shown in Fig. 25, although the amplitude becomes smaller. These results support that the assumption underlying the definition of $\kappa_{xy}$ remains valid over a broad range of temperatures and magnetic fields.
Figure 25: Semi-log plot of the position ($l$) dependence of the energy current amplitude for the pure Kitaev model at several values of the external magnetic field parallel to the $[111]$ direction at (a) $T\simeq 0.0373$, (b) $T\simeq 0.0745$, and (c) $T\simeq 0.1490$.
Appendix E Modulation of the magnetic state near the edges
In this Appendix, we discuss the magnetic state near the edges in the pure Kitaev model under the $[111]$ magnetic field. Due to the presence of open boundaries, the magnetic state near the edges can be different from the bulk magnetization. In Fig. 26, we plot the position dependence of the local magnetization along the $[111]$ direction ($M_{[111]}$) and $11\bar{2}]$ direction ($M_{[11\bar{2}]}$). Here, we define the magnetization at position $l$ as that of the top site in Fig. 1(b). In the center of the system ($l\sim 6$), $M_{[11\bar{2}]}$ almost vanishes and the magnetization is almost parallel to the $[111]$ direction, while spins near the edge $l=1$ exhibit a canting to the $11\bar{2}]$ direction. This spin canting could be attributed to the absence of the $z$-bond at the edge, suggesting that it is energetically favorable to have larger $x$ and $y$ components of the magnetization. It is worthy noting that, similarly to the energy current amplitude, the magnetization along the $11\bar{2}]$ direction also exhibits an almost exponential decay as a function of the position $l$ toward the center region, as shown in Fig. 26(b). This indicates that this magnetic modulation may contribute to the energy current localized at the edge.
Figure 26: The position ($l$) dependence of the local magnetization along the $[111]$ and $11\bar{2}]$ directions for the pure Kitaev model at several values of temperature under the magnetic field parallel to the $[111]$ direction with $h=0.04$ in (a) linear scale and (b) semi-log scale. Here, the local magnetization is defined at the top sites in each position ($l$) on the lattice in Fig. 1(b).